A particle of mass 2 kg is moving under the influence of a force F(x) = (3/vx) N. If the velocity of the particle at x = 2 m is v = 6 m/s, what is its speed at x = 7 m


1
Expert's answer
2021-10-18T11:08:32-0400

"a=\frac Fm=\frac 3{2\sqrt x}~\frac m{s^2},"

"v_1=\int a=3\sqrt x+C,"

"6=3\sqrt 2+C,"

"C=6-3\sqrt 2,"

"v_2=3\sqrt 7+6-3\sqrt 2\approx 2.3~\frac ms."